Friday 04/13/12

9:20 A – 5:00P, 8:30 P – 12:00A

  • miniprep Taf and Tub samples
  • Register with GeneWiz and send samples for sequencing
  • STORM imaging of en-Epc-H3 continues, focus lock lost early this morning when moving to top of new strip.  Pause feature of dave fails during reset.
  • Started new run in new sub-folder to avoid over-writing.  750 has strong background large features in image.  I suspect this is resin because 750 never gives background in my tissue normally.
  • washing out primary antibodies in oligo-STORM sample.
  • Tiling large regions with Steve on STORM2: select x and y at bottom and press ‘g’ to start tiling.  Note x is veritcal span and y is horizontal.  System takes 1 image in the middle and then scans from top to bottom building columns left to right.
  • Formatting 4TB WD external for RAID (and NTFS): WD icon on windows popup (same as for ejecting USB).   Select RAID 1.  Check serial number to be sure you are formatting the correct drive if formatting through Tuck / Cajal.
  •  New sample of en-EPc-H3 etched for 1.5 hrs in new (cold stored, anhydrous) sodium ethoxide solution is switching beautifully and completely etched.  No crazy background in 750.
  • transformation largely failed, no colonies, including positive control.  Repeating ligation with 20 uL volume and 10 uL of gel-extract.
  • STORM run 5p – 8P
  • O/N STORM run, also en-EPc-H3
This entry was posted in Summaries and tagged , . Bookmark the permalink.